OpenStudy (anonymous):

Evaluate the dot product of (-1,2) and (3,3).

OpenStudy (anonymous):

(-1,2) dot (3,3) = -1*3+2*3 = -3+6 = 3

OpenStudy (anonymous):

use the law of cosines to find the value of cos Round answer to two decimal places A. 1.14 B. 0.21 C. 0.85 D. 0.33

OpenStudy (precal):

well you are missing more information in order to use the laws of cosine